### v1

# Ultrastable nanophotonic microcavities via integrated thermometry

Abstract
Integrated photonic devices that can be co-packaged with electronics and can operate in real-world environments will enable many important applications, such as optical interconnects, quantum information processing, precision measurements, spectroscopy, and microwave generation. Significant progress has been made over the past two decades on increasing the functional complexity of photonic chips. However, the most critical challenge that remains is the lack of scalable techniques to overcome perturbations arising from environmental thermal noise and thermal crosstalk from co-packaged electronics and other photonic devices sharing the same substrate. Here, we propose and demonstrate a fully-integrated scheme to monitor and stabilize the temperature of a high-$Q$ microresonator in a Si-based chip. We show that when stabilized, the microresonator exhibits remarkable resilience against external thermal noise and can serve as a fully-integrated photonic frequency reference. By simply changing the temperature set-point, the cavity resonance frequency can be repeatably tuned without any hysteresis over several hours. We also show that the frequency of a distributed feedback (DFB) laser can be stabilized to this microresonator and realize a 48× reduction in its frequency drift, resulting in its center wavelength staying within ±0.5 pm of the mean over the duration of 50 hours in the presence of significant ambient fluctuations. This performance is superior to many commercial DFB systems and is highly suitable for use in data-communication systems. Finally, we demonstrate that the technique can be implemented to stabilize a soliton modelocked Kerr comb against significant ambient and crosstalk thermal noise, without the need for photodetection, paving the way for Kerr-comb-based photonic devices that can operate in the desired modelocked state indefinitely.
preprint: APS/123-QED
I Introduction
Optical frequency references are critically important tools for metrology [1, 2, 3], optical atomic clocks [4, 5, 6], precision spectroscopy [7, 8, 9], and telecommunication systems [10, 11, 12]. At wavelengths away from molecular, atomic, and nuclear transitions, stable optical cavities with narrow linewidths and low long-term drift are indispensable for laser frequency stabilization. A laser’s frequency can be stabilized by measuring its frequency detuning from the resonance of a stable narrow-linewidth cavity, and performing electronic or optical feedback correction. Within the bandwidth of the feedback mechanism, the cavity’s frequency stability can be transferred to the laser. To this end, cavities with large optical mode volumes offer exceptional fundamental frequency stability due to their low Brownian thermorefractive noise limit. For example, bulk Fabry-Perot (FP) cavities that utilize single-crystal Silicon and low-thermal-coefficient mirror coatings and housings operating at cryogenic temperatures have been reported to exhibit a fractional frequency stability of $10^{-17}$ over one second [13, 14]. Despite their exceptional stability, cryogenically-cooled bench-top systems are not well-suited for practical applications that require compactness, portability, and robustness. The rapid development of integrated low-loss dielectric photonic platforms over the past two decades offers the prospect of realizing stable reference microcavities in small form factors [15, 16]. However, the development of stable microcavities is challenging due to the enhanced sensitivity to fundamental and ambient temperature fluctuations that occur with small mode volumes, which in turn leads to significant drift in the cavity’s absolute resonance frequencies [17, 18].
High-$Q$ microcavities have proven to be excellent platforms for the generation of frequency combs, which have a number of applications such as spectroscopy [19, 20, 21], chip-scale multi-wavelength sources [22, 23], microwave generation [24, 25, 26], light detection and ranging (LiDAR) [27, 28], and optical interconnects for large-scale machine learning [29, 30]. Although monolithically-integrated microcavities offer superior mechanical stability than that of bulk cavities, a central challenge that remains is the lack of resource-efficient techniques to stabilize against thermal perturbations arising from ambient sources and thermal crosstalk from other thermally-tuned devices sharing the same substrate [31, 32]. For Kerr-comb applications, small temperature fluctuations can readily lead to pump detunings that fall outside of the modelocked regime of operation. Thermal crosstalk is a key bottleneck for neuromorphic computing applications [32] and densely-integrated co-packaged photonic input/output (I/O) modules [30], where multiple photonic devices on the same chip are typically simultaneously operated.
Passive methods to reduce Brownian thermal noise has been demonstrated previously by leveraging the inverse relationship with the mode volume, using a large-mode-volume integrated spiral cavity [15]. Manipulation of heat-flow, such as using undercuts [33] and trenches [34], have been shown to partially reduce the impact of external thermal perturbations, while adding significantly to fabrication cost and complexity. Active feedback using the photoconductive effect in silicon has been previously reported for stabilizing the temperature of microresonators [35, 36, 37]. Such a method, however, is inherently limited by the phase and intensity noise of the probing laser, and requires the use of highly-stable bench-top laser systems. Additionally, because the method relies upon the photoconductive effect to indirectly infer cavity temperature, its relevance is limited to semiconductor materials, and is not applicable to dielectric platforms that are commonly used in Kerr-comb-related applications. Laser frequency stabilization via dual-mode optical thermometry using a high-$Q$ microresonator has also been reported [38, 16]. However, it requires optical probing involving multiple phase modulators and cumbersome electronics, the excitation of two polarization or frequency modes, as well as exceptionally high quality factors for both two excited modes. All of these drawbacks make such a technique ill-suited for practical applications that require compactness and resource-efficient designs. Finally, for comb-based devices, active-feedback stabilization based on monitoring the optical output has been reported [39, 40]. However, the requirement of additional drop-ports, optical filters, photodiodes, and in many cases, free-space or fiber optics, significantly adds to the footprint and fabrication cost.
Here, we demonstrate a highly general approach, based on fully-integrated thermometry, to stabilize the temperature, and thereby the absolute resonance frequency, of a high-$Q$ microcavity in the presence of significant external noise over $24$ hours. A thin-film Platinum resistor is designed to exhibit a strong temperature-dependent electrical resistance arising from the metal’s intrinsic temperature coefficient of resistivity (TCR). Similar metallic structures have previously been extensively used as microheaters [41, 42, 43, 44]. We utilize such a resistor as an integrated resistance thermometer, enabling real-time monitoring of a microcavity’s temperature without the need for any external photodetection or other integrated non-linear electronic elements (e.g., diodes, transistors). The use of sputtered thin-films in our fabrication eliminates the need for processes requiring high-temperature and/or corrosive processing (e.g., electroplating). Moreover, Platinum being chemically resistant to basic chip cleaning reagents as well as atmospheric humidity leads to device longevity while ensuring long-term repeatability of our measurements with negligible variations. By engineering a dual-resistor scheme that ensures proper thermal stability of the microresonator core, we show that the absolute resonance frequency of a high-$Q$ silicon nitride (SiN) microcavity can be stabilized even in highly noisy environments to within tens of MHz of any desired value and can be repeatably tuned by simply adjusting an electrical voltage. The frequency noise of such a microcavity is shown to be impervious to significant environmental drift and thermal crosstalk, and is only limited by flicker noise in the control electronics. The resonance frequency of the stabilized microresonator exhibits bounded fluctuations over extended durations of time, in contrast to the unbounded random-walk behavior demonstrated by free-running microresonators and compact distributed feedback (DFB) lasers. By locking a DFB laser to the stabilized microresonator, we demonstrate that the microresonator can serve as a fully-integrated photonic frequency reference for applications that require long-term stability in a small form factor while operating in noisy environments, such as in integrated photonic transceivers for dense wavelength-division multiplexing (DWDM) applications. The locked laser exhibits a long-term stability of emission center wavelength of <±$<\pm 0.5$ pm over $50$ hours of measurement duration, significantly better than many commercial systems. Finally, we apply integrated thermometry to demonstrate a fully-thermally-stabilized Kerr frequency comb that remains stably modelocked in the presence of strong thermal crosstalk. We envision this work to enable the realization of highly robust and field-deployable linear and nonlinear photonic devices for classical and quantum applications, and integrated Kerr combs that can operate indefinitely.
II Integrated Thermometry
A conceptual illustration of integrated thermometry is provided in Fig. 1(a). A thin-film metallic resistor with a temperature-dependent resistivity, placed directly above a high-$Q$ microcavity, acts as an on-chip resistance thermometer. As a result of the low heat capacity of the thin-film resistor, small changes in heat flow can translate into large and observable changes in its temperature, and consequently in its electrical resistance. This enables monitoring changes in the microcavity temperature due to external noise, which is especially critical in practical scenarios. The current-voltage-resistance (I-V-R) characteristics of the metallic resistor, experimental measurements of which are shown in Fig. 1(b), indicate a deviation from the linear I-V trend, corresponding to a voltage-dependent resistance. This qualitative behavior is identical to a filament lamp, which exhibits lower (higher) resistance at lower (higher) voltages due to lower (higher) filament temperature arising from Joule heating. The near-quadratic dependence of resistance on applied voltage indicates a linear dependence on the power dissipated. Figure 1(c) shows a similar relationship between the thermometer’s resistance and the voltage applied across the second heater, demonstrating that temperature changes arising from external heat sources can be measured by simply monitoring the electrical resistance of the resistance thermometer.
Figure 1: a, Conceptual illustration of integrated thermometry applied to stabilize a high-Q monolithic microresonator against thermal noise arising from ambient heat sources and crosstalk from other thermally-tuned devices on the same chip. A thin-film metallic (platinum) resistor, typically used as a microheater, exhibits a temperature-dependent resistance due to the metal’s natural temperature-dependent resistivity. As a result of the low heat capacity of the thin-film resistor, small fluctuations in the heat flow in its vicinity lead to large changes in its temperature, enabling the real-time monitoring of on-chip temperature by simply measuring its electrical resistance. A second identical resistor is used as a heater to perform active stabilization purely using the resistance thermometer, thereby eliminating the need for optical probing for stabilization. Image for data center sourced from Ref. [45]. b, Current-voltage-resistance (I-V-R) characteristics of the resistance thermometer fabricated in this work. The I-V measurements (pale blue dots) deviate from the linear trend (dark blue line), and the resistance (yellow dots) exhibits a quadratic dependence (purple line) on the thermometer voltage. c, The measured thermometer resistance (lavender dots) also exhibits a quadratic dependence (green line) on the voltage applied across the (second) heater, indicating a linear dependence on power dissipated across the heater. d, Resonance frequency shift (green circles) of a high-Q microcavity of free-spectral range (FSR) 76 GHz, measured using a piezo-tuned tunable laser source, and measured change in thermometer resistance (gold line), for a sinusoidal perturbation applied via the heater. The cavity temperature and resonance frequency exhibit a strong anti-correlation. e, Measured resonance frequency shift plotted versus temperature measured by the thermometer, same data as in d. The linear fit provides an in-situ measurement of the thermo-optic coefficient of the fundamental spatial mode of the cavity waveguide.
The temperature measurement scheme can be utilized together with optical measurements to perform rapid in-situ measurements of the thermo-optic coefficient of the microcavity waveguide’s spatial mode at any desired wavelength. In order to demonstrate this, a small-amplitude 2-Hz sinusoidal voltage waveform is applied to the heater, while monitoring the electrical resistance of the thermometer. The free-spectral-range (FSR) of the SiN microresonator is 76 GHz, its loaded $Q$ is approximately $3\times 10^{6}$, and the resonances in the telecommunications C-band are critically coupled. The width of the waveguide is 2100 nm. A tunable laser source centered at 1560 nm is swept at 200 Hz, using its internal piezo-electric module, across a single microcavity resonance. By synchronously monitoring the transmitted optical output using a photodiode connected to a fast oscilloscope, the resonance frequency shift of the cavity is measured. Simultaneously, the on-chip temperature change is also monitored on the same oscilloscope by sending a constant low-noise DC current through the resistance thermometer and measuring the change in voltage across it. Figure 1(d) shows the measured resonance frequency shift (green circles; left y-axis) and the corresponding change in thermometer resistance (golden solid line; right y-axis), indicating a clear correlation between the two. The measured change in thermometer resistance is converted to a change in temperature using the following equation,
$R(T)=R_{0}(1+A\Delta T)$ (1)
where $R_{0}$ is the resistance at room temperature, $\Delta T$ is the change in temperature, and $A$ is the temperature coefficient of resistance (TCR). The coefficient $A$ was measured independently to be $A=0.00172$. The measurement procedure is detailed in Methods.
Figure 1(e) shows the data presented in Fig. 1(d), now with the resonance frequency shift plotted against the measured change in temperature. The slope of the linear fit is then used to extract the thermo-optic coefficient of the fundamental spatial mode of the waveguide, $dn_{eff}/dT=2.45\times 10^{-5}$ K-1. This value is also approximately equal to the thermo-optic coefficient $dn/dT$ of SiN due to the high spatial confinement offered by the fundamental mode of the 2100-nm waveguide, and is consistent with values published previously that employ different measurement techniques [46, 47, 48].
III Fully-Integrated Photonic Frequency Reference
The frequency stability of the microresonator is investigated, as depicted in Fig. 2(a), by piezo-scanning a probe laser across a cavity resonance at 200 Hz, and recording the resonance frequency position once every ten seconds. Absolute wavelength calibration is performed by beating the tunable laser with a low-drift fixed-wavelength reference laser on a fast photodiode, and simultaneously calibrating the reference laser’s wavelength using a benchtop wavemeter (Bristol Instruments 238A). This yields a calibrated absolute cavity resonance frequency, which is recorded continually once every 10 seconds over the course of 24 hours. The microcavity’s frequency stability is measured under the following three conditions: (i) ‘free-running’, in which the cavity is subject only to ambient fluctuations (such as change in laboratory temperature), (ii) ‘open-loop’, in which an additional perturbation is induced via a voltage summing circuit connected to the heater, in order to emulate real-world use, and (iii) ‘stabilized’, in which the microresonator’s temperature is actively stabilized. The strength of the perturbation added is denoted by the parameter $\xi$, which is defined as follows. For $\xi$=1, the root mean-squared (RMS) fluctuation of the cavity resonance frequency equals its short-term resonance full-width-at-half-maximum (FWHM) linewidth, which is 75 MHz for the microcavity used in this work. The cavity is stabilized by sensing the temperature in real-time by sending a low-noise electrical current through the resistance thermometer and measuring voltage changes across it, and by performing active feedback to the heater using a low-noise proportional-integral-derivative (PID) circuit. A schematic of the current source fabricated for this work and its noise performance are provided in Supplementary Information. The thermal design of the dual-resistor scheme described here is explained further in Methods.
Figure 2: Long-term stabilization of the absolute resonance frequency of a high-Q SiN microresonator using integrated thermometry. a, Simplified experimental schematic, depicting a tunable probe laser that is piezo-scanned across a microcavity resonance. Absolute frequency calibration of the scan is performed by producing a heterodyne note between the tunable laser and a stable reference laser, and simultaneously tracking the reference laser’s wavelength drift using a highly-stable benchtop wavelength meter. The heterodyne beat note as well as the cavity resonance line-shape are measured synchronously with the piezo sweep using an real-time oscilloscope connected to a computer interface for periodic data acquisition every 10 seconds. For $\xi$=1, the pseudorandom perturbation applied via the outer heater yields a root mean-squared (RMS) of the induced cavity resonance frequency fluctuation equals its short-term resonance full-width-at-half-maximum (FWHM) linewidth, which is 75 MHz for the microcavity fabricated in this work. b, Drift in the microcavity’s calibrated resonance frequency over the duration of 24 hours, for free-running ($\xi$=0), open-loop ($\xi$=7.55), and stabilized ($\xi$=7.55) operation. The free-running cavity exhibits a slow but strong drift in its resonance frequency due to ambient temperature drift in the lab, whereas the open-loop cavity additionally exhibits strong “fast” fluctuations due to the perturbation introduced. The stabilized cavity’s resonance frequency remains highly stable in the presence of very strong ambient and crosstalk-induced perturbations. c, Allan Deviation (ADEV) versus averaging time, computed using the data displayed in b. The stabilized cavity exhibits an ADEV that is well below the free-running and open-loop modes of operation. The shape of the ADEV curve of the stabilized cavity matches that of $1/f$ noise, indicating that performance is only limited by control electronics. d, Histogram plots depicting the fraction of time over 24 hours (y-axis) that the microcavity’s absolute resonance frequency drift lies within a given frequency bin with a width of 5 MHz.
Figure 2(b) displays the measured resonance frequency drift of the microresonator for the three cases described above over a period of 24 hours. In the absence of any stabilization, ambient noise alone causes a significant resonance drift, even in temperature-stabilized laboratory environments. This is shown by the slow cyclical pattern in the ‘free-running cavity’ data points, with an standard deviation of resonance frequency variation of $\approx 341$ MHz. Additional perturbation, such as from thermal crosstalk in practical applications, adds to the drift in the form of faster-timescale fluctuations, as evidenced by the ‘open-loop cavity’ data points, exhibiting a standard deviation of resonance frequency variation of $\approx 470$ MHz. The Supplementary Information provides the corresponding real-time temperatures measured by the resistance thermometer for the free-running and open-loop cases, which shows excellent correlation with the resonance frequency shift. Finally, the stabilized cavity exhibits a highly stable resonance frequency with a considerably lower drift even in the presence of significant added perturbation ($\xi=7.55$). The standard deviation of the measured resonance frequency variation of the stabilized microresonator is $\approx 13.7$ MHz, corresponding to a 34×34$\times$34 × reduction compared to the open-loop scenario. The time-domain measurement data is better visualized as histograms, shown in Fig. 2(d). The bin-size along the x-axis is 5 MHz. Remarkably, even in the absence of additional perturbations, i.e., in the free-running case, the cavity resonance frequency histogram does not consist of a single, well-defined narrow peak.
Figure 2(c) shows the numerically-computed Allan deviation (ADEV). As expected, the stabilized cavity offers a significantly lower ADEV compared to the free-running and open-loop cavities. The shape of the ADEV of the stabilized cavity closely matches that of pure $1/f$ noise, indicating that unbounded random walk resonance frequency drifts arising from environmental and crosstalk perturbations are effectively canceled out by the stabilization method employed and that the residual fluctuations are bounded, limited only by noise in the control electronics used in this demonstration. This is evidenced by the flat shape of the ADEV versus averaging time, compared to the free-running and open-loop cavities, which exhibit non-stationary unbounded random walks, shown by the increasing ADEV at large averaging times. The stabilized microresonator also offers significantly higher long-term stability compared to temperature-stabilized DFB lasers, which are commonly used in chip-based telecommunications applications [49, 30], making it a highly practical fully-integrated frequency reference cavity.
Finally, long-term repeatability of the microcavity’s tunable absolute resonance wavelength is assessed as follows. To assess the long-term repeatability of a thermally-tuned cavity only subject to ambient thermal noise (i.e., $\xi=0$), the voltage applied to the outer heater is swept in discrete steps so as to trace the logo of Columbia University, and the microcavity’s absolute resonance wavelength is synchronously measured using the scanning probe laser setup described in Fig. 2(a). This measurement is repeated in the presence of additional perturbations to mimic external thermal fluctuations in practical photonic devices ($\xi=3.77$). No feedback is applied in the first two configurations. Finally, the feedback loop is closed, and thermal tuning is performed by simply changing the set-point of the PID loop. The results are shown in Fig. 3.
Figure 3: Demonstration of tunability and long-term repeatability of the microresonator’s absolute resonance wavelength. The resonance frequency is programmed to trace a pre-defined shape by blind tuning of the power supplied to the outer heater in the presence of ambient noise alone ($\xi$=0; shown in a), in the presence of additional perturbation ($\xi$=3.77; shown in b), and by tuning of the set-point for temperature stabilization in the presence of ambient noise as well as additional perturbation ($\xi$=3.77; shown in c). The measurements for heater-only tuning show that the microresonator’s resonance wavelength exhibits a large drift over the timescale of only a few hours, even for the same pre-programmed heater voltage. The stabilized tuning, on the other hand, yields a highly repeatable performance even in the presence of strong ambient and external noise.
When tuned with the heater alone, i.e., the most common way of tuning microresonators, applying the same voltage does not lead to the same resonance wavelength over the measurement duration of 10 hours, as evidenced by the drift in the patterns in Figs. 3(a) and (b). Similar to the results presented in Fig. 2, additional perturbations lead to “short-term” fluctuations superimposed on top of the long-term drift caused by environmental drift. However, in the stabilized case in which the set-point of the feedback loop is tuned, active feedback based on the resistance thermometer’s “reading” cancels out ambient drift as well as the perturbation-induced fluctuations, and the microcavity resonance wavelength exhibits significantly better repeatability.
The long-term frequency stability of the thermometrically stabilized microresonator is demonstrated further by stabilizing the frequency of a laser. A DFB laser is locked to a microcavity resonance using the Pound-Drever-Hall (PDH) method, and its emission center wavelength is recorded every 10 seconds using a bench-top wavelength meter. We employ the PDH method in this experiment due to its insensitivity to drift in the input fiber-to-chip optical coupling. In a fully-integrated system, however, the frequency stability of the microresonator is easily transferred to an integrated laser via mechanisms such as self-injection locking [50]. The results of the frequency stabilization demonstration are shown in Fig. 4.
Figure 4(a) shows the measured drift in laser center wavelength for the free-running and PDH-locked DFB lasers, in blue and maroon respectively. The mean emission wavelength for the two cases are approximately 1551.2 nm. The free-running DFB experiences significant wavelength drift arising from ambient temperature fluctuations alone, amounting to a standard deviation of $11.5$ pm (i.e., $1.43$ GHz) over the span of $42$ hours of measurement time. In comparison, the locked DFB emits at a highly stable center wavelength, with a peak-to-peak variation that lies within ±$\pm 0.5$ pm of the mean over $50$ hours of measurement time, and exhibits a standard deviation of $0.24$ pm. The stability enhancement factor is 48$\approx$ 48$\times$ 48 ×. The long-term stability measured here is significantly superior to many commercially-available DFB laser system, and makes the scheme presented here ideal for the realization of practical photonic transceivers for dense wavelength division multiplexing (DWDM) applications.
Figure 4(b) shows that the computed ADEV of the locked DFB is significantly lower ADEV than the free-running DFB, and is comparable to the ADEV of the stabilized microresonator displayed in Fig. 2. By using a large integral gain in the feedback servo of the PDH, a slightly lower ADEV is observed at smaller averaging times. The slightly higher ADEV at larger averaging times is attributed to insufficient gain of the PDH feedback servo at very low frequencies.
Figure 4: Demonstration of long-term laser frequency stabilization to the microresonator. A DFB laser is locked to the resonance of the stabilized microresonator using the Pound-Drever-Hall (PDH) method, and its emission center wavelength is continually measured once every 10 seconds using a bench-top wavelength meter. a, Time-domain comparison of center wavelength drift of free-running and PDH-locked configurations of a DFB laser. The locked laser’s center wavelength remains within a ±$\pm 0.5$ pm window of the mean, indicated by the dashed lines, for 50 hours. The free-running laser is only subject to ambient temperature drift in a temperature-stabilized laboratory environment. b, Comparison of measured Allan Deviation (ADEV) of the free-running and locked configurations of the DFB laser. The locked laser exhibits slightly lower ADEV than the stabilized microresonator at lower averaging times due to the strong integral gain of the PDH feedback servo. The slightly higher ADEV at larger averaging times is attributed to insufficient gain of the PDH feedback servo at very low frequencies.
IV Fully-thermally-stabilized Kerr Comb
Finally, we show that the fully-integrated thermometry approach described in this work can stabilize a modelocked Kerr comb, which is typically highly sensitive to perturbations, for long periods. The experimental schematic is shown in Fig. 5(a). A pump laser at approximately 1550 nm is amplified before coupling into the microcavity device. The off-resonance optical power output is 150-mW. The low-noise current source connected to the resistance thermometer is modulated with a burst waveform to generate a single-soliton Kerr comb, shown in blue in the optical spectrum analyzer (OSA) measurements in Fig. 5(b). A portion of the out-coupled Kerr comb is sent through a fiber-optic notch filter centered at the pump wavelength, and the transmitted output, shown in red in Figure 5(b), is used to track the comb power using a photodiode. Phase-noise characterization is performed by generating a microwave signal at the repetition rate of the comb (76 GHz) via a high-speed photodetector (bandwidth 100 GHz), and measuring the microwave phase noise using a phase noise analyzer (PNA; Rohde & Schwarz ®  FSWP50). Since the microwave frequency (76 GHz) is beyond the specified bandwidth (50 GHz) of the PNA used in this work, an 8:1 harmonic mixer (Rohde & Schwarz ®  RPG FS-Z110) is employed to down-convert the microwave frequency to within the PNA’s bandwidth.
The stability of the Kerr comb is studied by applying a perturbation (white-noise voltage, bandwidth $=20$ Hz) via the outer on-chip heater and measuring the real-time on-chip temperature and the power and phase noise of the comb. For a noise strength of $\xi=1.51$, without any feedback correction, the modelocked Kerr comb transitions to a chaotic comb state almost immediately, as shown in Fig. 5(c), where the top plot shows the applied thermal perturbation voltage signal in the time-domain. The thermometer voltage plot, for a constant low-noise current sent through the resistance thermometer, shows a sudden increase in the cavity temperature, indicating a sudden jump to a state with higher intracavity power. This is explained by the bottom plot, which shows that the comb power (which is measured by filtering out the pump line) is no longer stable and exhibits random fluctuations driven by the perturbation voltage. These fluctuations are also seen in the thermometer voltage following the jump. The smaller amplitude of fluctuations observed by the thermometer is a result of an increase (decrease) in temperature at the thermometer due the perturbation and the consequent secondary effect, namely decrease (increase) in temperature at the thermometer due to decrease (increase) in intracavity optical power and self-heating. This is a result of the inverse relationship between cavity temperature and intracavity power for the effectively blue-detuned pump of the chaotic comb state, which does not apply for the modelocked comb state, since the pump is effectively red-detuned, and the resistance thermometer therefore remains sensitive to external perturbations, making it an efficient tool in stabilizing modelocked comb states.
Figure 5: Fully-thermally stabilized Kerr optical frequency comb. a, Simplified experimental schematic, depicting the comb generation and characterization setup. A soliton Kerr comb is generated in the microresonator by fast-tuning the current supplied to the resistance thermometer. The output comb is measured using an optical spectrum analyzer (OSA), and its frequency noise is measured using a high-speed photodetector connected to a phase-noise analyzer (PNA) via an 8:1 harmonic mixer (HM). The comb power is tracked using a slow photodiode (PD) after passing the output light through an optical notch filter (ONF). Comb stability in the presence of external noise is studied by applying a perturbation across the outer heater after generating the comb. b, Generated single-soliton and filtered output spectra, measured using an OSA. c and d, Perturbation voltage (top), measured temperature (middle), and output comb power (bottom), for open-loop and stabilization operations, for noise strengths of $\xi$=1.51 and $\xi$=7.55 respectively. Without stabilization, the comb jumps from a modelocked state to a chaotic regime almost instantly after the perturbation amplitude is increased. When stabilized, the comb remains modelocked even in the presence of much stronger noise. e, Measured single-sideband (SSB) frequency noise of the microwave generated by detecting the comb using a very-high-speed PD, for open-loop operation for a noise strength that is not weak enough to completely destabilize the comb ($\xi$=0.75; yellow line) and stabilized operation for a high noise strength ($\xi$=7.55; purple line). The microwave generated in the stabilized case has considerably lower noise than the open-loop case despite the much stronger noise.
The stability of the closed-loop system is studied by repeating the same measurement sequence as above. However, in this case, the noise strength is increased five-fold to $\xi=7.55$. As seen in Fig. 5(d), the PID correction voltage cancels out the applied perturbation, leading to a stable temperature and Kerr comb power even in the presence of strong perturbations. Whereas the Kerr comb is pushed out of its existence range immediately in the open-loop ($\xi=1.51$) scenario, closed-loop operation was observed to support modelocking for over 20 minutes, limited only by the drift in the input fiber-to-chip optical coupling, paving the way for Kerr-comb-based photonic devices that can operate in the desired modelocked state indefinitely.
Lastly, Fig. 5(e) shows the comparison of the microwave single-sideband (SSB) frequency noise. For the ‘open-loop’ scenario, the noise strength is reduced to $\xi=0.75$ to keep the comb in the modelocked regime of operation. For the ‘stabilized’ scenario, the noise strength is maintained at $\xi=7.55$. As the yellow and purple solid curves show, the frequency noise of the microwave generated by the stabilized cavity is over 20 dB below that of the microwave generated by the open-loop cavity. At lower offset frequencies (i.e., over longer timescales), the stabilized cavity’s microwave exhibits a lower noise than the free-running cavity’s microwave since ambient drift is more prominent on longer timescales.
V Conclusion
In conclusion, we demonstrate a powerful new approach based on fully-integrated thermometry to stabilize the absolute resonance frequency of high-$Q$ silicon nitride microcavities without requiring more complicated monitoring systems such as photodetection. The thermometric stabilization mitigates environmental drift as well as externally-induced perturbations effectively, and the Allan deviation of the stabilized microcavity’s resonance frequency remains nearly flat versus averaging time, in contrast to the unbounded random walks exhibited by the free-running and open-loop cases, evidenced by the increasing Allan deviation at large averaging times. By locking a DFB laser to the stabilized microresonator, we demonstrate that the center-wavelength stability of the laser is <±$<\pm 0.5$ pm from the mean over the measurement duration of $50$ continuous hours, which is a 48×48$\times$48 × improvement compared to the free-running laser. This work paves the way for highly compact, fully-integrated and field-deployable frequency references for telecommunications applications.
When pumped with a high-power laser source to generate a Kerr comb, the thermometry-based stabilization technique described in this work makes the comb significantly more stable even in the presence of strong perturbations. The fully-thermally-stabilized comb demonstrated here offers a path toward comb-based high-$Q$ photonic devices that can operate indefinitely even in noisy environments.
Although the demonstration in this work was limited to the SiN platform, this technique is readily generalizable to other photonic materials. Notably, previously-demonstrated on-chip temperature-sensing techniques, such as ones leveraging the photoconductive effect, require the presence of significant optical power in the cavity, and are limited to semiconductor platforms. The resistance-thermometer-based scheme here poses no such requirement. Finally, removing the requirement of additional drop-ports and photodetection for stabilization can reduce the design complexity and fabrication costs of practical photonic devices quite significantly.
VI Acknowledgments
The authors wish to acknowledge Kaylx Jang and Michael Cullen from the group of Prof. Keren Bergman for providing the photograph of a co-packaged silicon photonics device, previously reported in Ref. [51], included in Fig. 1. The authors thank Vivian Zhou from the group of Prof. Michal Lipson for assistance in setting up the vacuum system used in the temperature coefficient of resistance measurements.
VII Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
VIII Funding
The authors acknowledge support from the Army Research Office under Cooperative Agreement Number W911NF-24-2-0204, the Air Force Office of Scientific Research (FA9550-20-1-0297), and the Defense Advanced Research Projects Agency (HR0011-19-2-0014).
Ultrastable nanophotonic microcavities via integrated thermometry: Methods
I Thermal Design for Integrated Thermometry
Figure M1: Thermal design for integrated thermometry. a, Numerically simulated steady-state temperature profile for the scheme described in this work. The heater power is set to $100$ mW for the purposes of illustration. The initial temperature of the chip is $293$ K. b, Heater-thermometer spacing design trade-off, showing the core-thermometer temperature differential (left y-axis, in blue), and the heater power required for core heating of $0.1$ K (right y-axis, in dark red), versus heater-thermometer spacing.
Proper thermal design is essential for realizing an on-chip temperature-stabilization scheme that is effective at stabilizing the resonance frequency of the high-$Q$ microresonator. Specifically, the heater must be placed at a radial separation from the microresonator such that the any thermal correction it applies leads to equal temperature changes at the resistance thermometer as well as the microresonator core. The right heater-thermometer spacing is determined prior to device fabrication by solving the following heat transfer equation in steady-state using the COMSOL Multiphysics® software:
$\rho C_{p}\frac{\partial T}{\partial t}+\nabla\cdot\bm{q}=Q+Q_{ted}$ (M1)
where $\rho$ is the material density, $C_{p}$ is the specific heat capacity at constant stress, $T$ represents the temperature, $q=-k\nabla T$ is the heat flux by conduction, $k$ is the thermal conductivity, $Q$ contains additional heat sources, and $Q_{ted}$ accounts for thermo-elastic damping in the material. All material thermal properties were set by the standard COMSOL material database for the cross-section of our devices. Figure M1(a) displays the cross-section of our device, and the steady-state spatial temperature profile of the stationary thermal simulation. The heat power is set to $100$ mW for the purposes of illustration. The initial temperature of the entire chip, and the boundary condition temperature for the bottom face of the chip, are set to $293$ K. Here, we denote the resulting steady-state temperature change at the center of the SiN core and at the center of the resistance thermometer as $\delta$Tcore and $\delta$Ttherm., respectively.
Figure M1(b) shows the temperature differential $\delta$Tcore - $\delta$Ttherm. (left y-axis, in blue) as well as the heater power required to heat up the core (right y-axis, in dark red), which is a measure of heater efficiency, versus the heater-thermometer spacing. For small spacing, the heater-thermometer and heater-core separations are considerably different, leading to a large temperature differential. However, the close proximity allows for greater heater efficiency. For large spacing, the temperature differential drops, as the two separations are almost equal, while the heater efficiency worsens. The design employed in this work utilizes a $20$ $\mu$m heater-thermometer spacing, which leads to a very low temperature differential of $<5$ mK, making it effective in stabilizing the resonance frequency of the microcavity. The design principles described here can be readily applied to other integrated material platforms.
II Device Fabrication
The on-chip resistance thermometer demonstrated in this work is similar to devices that have been extensively used as resistive microheaters [41, 42]. It consists of a Platinum thin-film resistor deposited on top of the oxide cladding layer. The resistors are fabricated by first patterning on top oxide layer, a stack of lift-off resist and positive photo-resist. Blanket Ti-Pt films are sputter deposited at room temperature, followed by metal lift-off to create resistors. A thin layer ($\approx 5\ \textrm{nm}$) of Titanium (Ti) is first deposited before deposition of Pt ($\approx 100\ \textrm{nm}$), to promote adhesion of Pt films. The sputtered metal adheres to and remains on the exposed oxide surface while metal on the photo-resist is stripped away in a resist stripping solution. The use of sputtered thin-films in our fabrication eliminates the need for processes requiring high-temperature and/or corrosive processing (e.g., electroplating).
III Measurement of Temperature Coefficient of Resistance (TCR)
Figure M2: Measurement of temperature coefficient of resistance of the integrated Platinum thermometer. a, Experimental schematic depicting a test chip, containing a platinum resistor (referred to here as device under test (DUT)) fabricated on top of a SiN microresonator, placed on a temperature-stabilized hot plate chuck inside a vacuum chamber. Its electrical resistance is probed as a function of temperature by performing current-voltage characterization using a Keithley $2400$ series SourceMeter®. b, Measured electrical resistance of DUT versus temperature (dark-blue filled circles) plotted alongside a linear fit (magenta solid line).
The integrated Platinum resistance thermometer, which allows the real-time measurement of the microresonator as described in the Main Text, is calibrated as shown in Fig. M2(a). A test device is placed on a temperature-controlled chuck using a thermally-conductive adhesive in order to ensure proper thermal contact. The setup is placed and operated inside a vacuum probe station in order to minimize unwanted convective cooling. Two probes, connected to a Keithley $2400$ series SourceMeter®, are brought in electrical contact with the two ends of the Pt resistor under test (labeled as device under test (DUT) in Fig. M2). At a given temperature of the chuck, the resistance $R$ of the DUT is measured by performing a current-voltage characterization. The voltage drop across the two contact pads is measured at various source currents, and a linear fit of the resulting data is obtained to extract $R$. This method is further explained in a prior manuscript [52]. The source current values used are low enough ($<1$ mA) to avoid any self-heating of the resistor. This process is repeated for a range of chuck temperatures $T$. Figure M2(b) displays the results. Over the range of temperatures probed on this experiment, the resistance exhibits a linear trend. The temperature-coefficient of resistance is extracted by fitting the experimental data to Eq. 1, and is determined to be $A=0.00172$. We note that DUT described here has a shorter fabricated length compared to the resistance thermometer described in the Main Text, leading to the lower base resistance.

### v2

@@
-Ultrastable nanophotonic microcavities via integrated thermometry
+Frequency-stable nanophotonic microcavities via integrated thermometry
 Abstract
-Integrated photonic devices that can be co-packaged with electronics and can operate in real-world environments will enable many important applications, such as optical interconnects, quantum information processing, precision measurements, spectroscopy, and microwave generation.
+Field-deployable integrated photonic devices co-packaged with electronics will enable important applications such as optical interconnects, quantum information processing, precision measurements, spectroscopy, and microwave generation.
 Significant progress has been made over the past two decades on increasing the functional complexity of photonic chips.
-However, the most critical challenge that remains is the lack of scalable techniques to overcome perturbations arising from environmental thermal noise and thermal crosstalk from co-packaged electronics and other photonic devices sharing the same substrate.
-Here, we propose and demonstrate a fully-integrated scheme to monitor and stabilize the temperature of a high-$Q$ microresonator in a Si-based chip.
-We show that when stabilized, the microresonator exhibits remarkable resilience against external thermal noise and can serve as a fully-integrated photonic frequency reference.
-By simply changing the temperature set-point, the cavity resonance frequency can be repeatably tuned without any hysteresis over several hours.
-We also show that the frequency of a distributed feedback (DFB) laser can be stabilized to this microresonator and realize a 48× reduction in its frequency drift, resulting in its center wavelength staying within ±0.5 pm of the mean over the duration of 50 hours in the presence of significant ambient fluctuations.
-This performance is superior to many commercial DFB systems and is highly suitable for use in data-communication systems.
-Finally, we demonstrate that the technique can be implemented to stabilize a soliton modelocked Kerr comb against significant ambient and crosstalk thermal noise, without the need for photodetection, paving the way for Kerr-comb-based photonic devices that can operate in the desired modelocked state indefinitely.
+However, a critical challenge that remains is the lack of scalable techniques to overcome thermal perturbations arising from the environment and co-packaged electronics.
+Here, we demonstrate a fully integrated scheme to monitor and stabilize the temperature of a high-$Q$ microresonator on a Si-based chip, which can serve as a photonic frequency reference.
+Our approach relies on a thin-film metallic resistor placed directly above the microcavity, acting as an integrated resistance thermometer, enabling unique mapping of the cavity’s absolute resonance wavelength to the thermometer’s electrical resistance.
+Following a one-time calibration, the microresonator can be accurately and repeatably tuned to any desired absolute resonance wavelength using thermometry alone with a root-mean squared wavelength error of <0.8<0.8 pm over a timespan of days.
+We frequency-lock a distributed feedback (DFB) laser to the microresonator and demonstrate a 48× reduction in its frequency drift, resulting in its center wavelength staying within ±0.5 pm of the mean over the duration of 50 hours in the presence of significant ambient fluctuations, outperforming many commercial DFB and wavelength-locker-based laser systems.
+Finally, we stabilize a soliton mode-locked Kerr comb without the need for photodetection, paving the way for Kerr-comb-based photonic devices that can potentially operate in the desired mode-locked state indefinitely.
 preprint: APS/123-QED
-I Introduction
 Optical frequency references are critically important tools for metrology [1, 2, 3], optical atomic clocks [4, 5, 6], precision spectroscopy [7, 8, 9], and telecommunication systems [10, 11, 12].
-At wavelengths away from molecular, atomic, and nuclear transitions, stable optical cavities with narrow linewidths and low long-term drift are indispensable for laser frequency stabilization.
-A laser’s frequency can be stabilized by measuring its frequency detuning from the resonance of a stable narrow-linewidth cavity, and performing electronic or optical feedback correction.
-Within the bandwidth of the feedback mechanism, the cavity’s frequency stability can be transferred to the laser.
-To this end, cavities with large optical mode volumes offer exceptional fundamental frequency stability due to their low Brownian thermorefractive noise limit.
-For example, bulk Fabry-Perot (FP) cavities that utilize single-crystal Silicon and low-thermal-coefficient mirror coatings and housings operating at cryogenic temperatures have been reported to exhibit a fractional frequency stability of $10^{-17}$ over one second [13, 14].
-Despite their exceptional stability, cryogenically-cooled bench-top systems are not well-suited for practical applications that require compactness, portability, and robustness.
-The rapid development of integrated low-loss dielectric photonic platforms over the past two decades offers the prospect of realizing stable reference microcavities in small form factors [15, 16].
-However, the development of stable microcavities is challenging due to the enhanced sensitivity to fundamental and ambient temperature fluctuations that occur with small mode volumes, which in turn leads to significant drift in the cavity’s absolute resonance frequencies [17, 18].
-High-$Q$ microcavities have proven to be excellent platforms for the generation of frequency combs, which have a number of applications such as spectroscopy [19, 20, 21], chip-scale multi-wavelength sources [22, 23], microwave generation [24, 25, 26], light detection and ranging (LiDAR) [27, 28], and optical interconnects for large-scale machine learning [29, 30].
-Although monolithically-integrated microcavities offer superior mechanical stability than that of bulk cavities, a central challenge that remains is the lack of resource-efficient techniques to stabilize against thermal perturbations arising from ambient sources and thermal crosstalk from other thermally-tuned devices sharing the same substrate [31, 32].
-For Kerr-comb applications, small temperature fluctuations can readily lead to pump detunings that fall outside of the modelocked regime of operation.
-Thermal crosstalk is a key bottleneck for neuromorphic computing applications [32] and densely-integrated co-packaged photonic input/output (I/O) modules [30], where multiple photonic devices on the same chip are typically simultaneously operated.
-Passive methods to reduce Brownian thermal noise has been demonstrated previously by leveraging the inverse relationship with the mode volume, using a large-mode-volume integrated spiral cavity [15].
-Manipulation of heat-flow, such as using undercuts [33] and trenches [34], have been shown to partially reduce the impact of external thermal perturbations, while adding significantly to fabrication cost and complexity.
-Active feedback using the photoconductive effect in silicon has been previously reported for stabilizing the temperature of microresonators [35, 36, 37].
-Such a method, however, is inherently limited by the phase and intensity noise of the probing laser, and requires the use of highly-stable bench-top laser systems.
-Additionally, because the method relies upon the photoconductive effect to indirectly infer cavity temperature, its relevance is limited to semiconductor materials, and is not applicable to dielectric platforms that are commonly used in Kerr-comb-related applications.
+At wavelengths away from molecular, atomic, and nuclear transitions, optical cavities with narrow linewidths and low long-term drift are indispensable for laser frequency stabilization.
+Cavities with large optical mode volumes and low Brownian thermorefractive noise (TRN) offer the best frequency stability, such as bulk single-crystal Silicon Fabry-Perot (FP) cavities with a reported fractional frequency instability of $10^{-17}$ over one second [13, 14].
+For many applications that require portability and robustness, such as in telecommunication systems, cryogenically-cooled bulk cavities are not well-suited, necessitating the use of compact reference and laser cavities that trade off size, weight, and power consumption (SWaP) for frequency stability that is typically several orders of magnitude worse compared to state-of-the-art bulk systems.
+In addition to the plethora of commercially available temperature-stabilized distributed feedback (DFB) and distributed Bragg reflector (DBR) lasers, several compact frequency stabilization technologies have been demonstrated in recent years, including integrated optical frequency discriminators [15, 16], vapor-cell-integrated photonic chips [17], compact wavelength lockers [18], and fiber interferometers [19].
+The rapid development of low-loss dielectric photonic platforms over the past two decades offers the prospect of realizing stable reference cavities in small form factors [20, 21].
+In recent years, microcavities with large quality factors ($Q$s) have been shown to enable short-term frequency instability approaching $10^{-14}$ [22, 23].
+However, achieving long-term stability remains challenging due to the enhanced sensitivity to fundamental and ambient temperature fluctuations [24, 25].
+High-$Q$ microcavities are also excellent platforms for the generation of frequency combs, which have a number of applications such as spectroscopy [26, 27, 28], chip-scale multi-wavelength sources [29, 30], microwave generation [31, 32, 33], light detection and ranging (LiDAR) [34, 35], and optical interconnects for large-scale machine learning [36, 37].
+A central challenge that remains to be addressed in integrated photonics broadly is the lack of resource-efficient techniques to stabilize against thermal perturbations arising from ambient sources, co-packaged electronics, and thermal crosstalk [38, 39].
+For Kerr-comb applications, small temperature fluctuations can readily lead to pump detunings that lie outside of the modelocked regime of operation.
+Thermal crosstalk is a key bottleneck for neuromorphic computing applications [39] and densely integrated co-packaged photonic input/output (I/O) modules that require simultaneous tuning of multiple on-chip devices [37].
+Passive methods to mitigate thermal perturbations have been demonstrated previously, such as large mode volumes [20], and manipulating heat flow using undercuts [40] and trenches [41].
+However, passive methods only partially reduce the impact of thermal perturbations while adding significantly to fabrication cost and complexity.
+Active stabilization of silicon microresonators, based on the photoconductive effect, to a reference laser has been shown [42, 43, 44].
+However, such a method is inherently limited by the reference laser’s stability and is not applicable to high-$Q$ dielectric platforms that are commonly used for realizing Kerr combs.
 Laser frequency stabilization via dual-mode optical thermometry using a high-$Q$ microresonator has also been reported [45, 21].
-However, it requires optical probing involving multiple phase modulators and cumbersome electronics, the excitation of two polarization or frequency modes, as well as exceptionally high quality factors for both two excited modes.
-All of these drawbacks make such a technique ill-suited for practical applications that require compactness and resource-efficient designs.
-Finally, for comb-based devices, active-feedback stabilization based on monitoring the optical output has been reported [39, 40].
-However, the requirement of additional drop-ports, optical filters, photodiodes, and in many cases, free-space or fiber optics, significantly adds to the footprint and fabrication cost.
-Here, we demonstrate a highly general approach, based on fully-integrated thermometry, to stabilize the temperature, and thereby the absolute resonance frequency, of a high-$Q$ microcavity in the presence of significant external noise over $24$ hours.
-A thin-film Platinum resistor is designed to exhibit a strong temperature-dependent electrical resistance arising from the metal’s intrinsic temperature coefficient of resistivity (TCR).
+However, it requires multiple phase modulators, cumbersome electronics, and the excitation of modes with proximate resonance frequencies and very high $Q$s.
+For comb-based devices, while active feedback using the optical output has been reported [46, 47], the requirement of additional drop-ports, optical filters, photodiodes, and free-space/fiber optics significantly adds to footprint and fabrication cost.
+Here, we demonstrate a highly general approach, based on fully integrated thermometry, to stabilize the temperature of a high-$Q$ microcavity in the presence of significant external perturbations.
+A thin-film Platinum resistor is designed to exhibit a strong temperature-dependent electrical resistance and is used as an integrated resistance thermometer to measure a microcavity’s temperature without the need for any photodetection or other integrated non-linear electronic elements (e.g., diodes, transistors).
 Similar metallic structures have previously been extensively used as microheaters [48, 49, 50, 51].
-We utilize such a resistor as an integrated resistance thermometer, enabling real-time monitoring of a microcavity’s temperature without the need for any external photodetection or other integrated non-linear electronic elements (e.g., diodes, transistors).
-The use of sputtered thin-films in our fabrication eliminates the need for processes requiring high-temperature and/or corrosive processing (e.g., electroplating).
-Moreover, Platinum being chemically resistant to basic chip cleaning reagents as well as atmospheric humidity leads to device longevity while ensuring long-term repeatability of our measurements with negligible variations.
-By engineering a dual-resistor scheme that ensures proper thermal stability of the microresonator core, we show that the absolute resonance frequency of a high-$Q$ silicon nitride (SiN) microcavity can be stabilized even in highly noisy environments to within tens of MHz of any desired value and can be repeatably tuned by simply adjusting an electrical voltage.
-The frequency noise of such a microcavity is shown to be impervious to significant environmental drift and thermal crosstalk, and is only limited by flicker noise in the control electronics.
-The resonance frequency of the stabilized microresonator exhibits bounded fluctuations over extended durations of time, in contrast to the unbounded random-walk behavior demonstrated by free-running microresonators and compact distributed feedback (DFB) lasers.
-By locking a DFB laser to the stabilized microresonator, we demonstrate that the microresonator can serve as a fully-integrated photonic frequency reference for applications that require long-term stability in a small form factor while operating in noisy environments, such as in integrated photonic transceivers for dense wavelength-division multiplexing (DWDM) applications.
-The locked laser exhibits a long-term stability of emission center wavelength of <±$<\pm 0.5$ pm over $50$ hours of measurement duration, significantly better than many commercial systems.
-Finally, we apply integrated thermometry to demonstrate a fully-thermally-stabilized Kerr frequency comb that remains stably modelocked in the presence of strong thermal crosstalk.
-We envision this work to enable the realization of highly robust and field-deployable linear and nonlinear photonic devices for classical and quantum applications, and integrated Kerr combs that can operate indefinitely.
-II Integrated Thermometry
-A conceptual illustration of integrated thermometry is provided in Fig.
-1(a).
-A thin-film metallic resistor with a temperature-dependent resistivity, placed directly above a high-$Q$ microcavity, acts as an on-chip resistance thermometer.
-As a result of the low heat capacity of the thin-film resistor, small changes in heat flow can translate into large and observable changes in its temperature, and consequently in its electrical resistance.
-This enables monitoring changes in the microcavity temperature due to external noise, which is especially critical in practical scenarios.
-The current-voltage-resistance (I-V-R) characteristics of the metallic resistor, experimental measurements of which are shown in Fig.
-1(b), indicate a deviation from the linear I-V trend, corresponding to a voltage-dependent resistance.
-This qualitative behavior is identical to a filament lamp, which exhibits lower (higher) resistance at lower (higher) voltages due to lower (higher) filament temperature arising from Joule heating.
-The near-quadratic dependence of resistance on applied voltage indicates a linear dependence on the power dissipated.
-Figure 1(c) shows a similar relationship between the thermometer’s resistance and the voltage applied across the second heater, demonstrating that temperature changes arising from external heat sources can be measured by simply monitoring the electrical resistance of the resistance thermometer.
-Figure 1: a, Conceptual illustration of integrated thermometry applied to stabilize a high-Q monolithic microresonator against thermal noise arising from ambient heat sources and crosstalk from other thermally-tuned devices on the same chip.
+Platinum’s chemical stability against basic chip cleaning reagents and atmospheric humidity ensures device longevity and long-term repeatability of our measurements.
+By engineering a dual-resistor scheme that achieves proper thermal stability of the microresonator waveguide core, we show that the absolute resonance frequency of an unshielded and unpackaged high-$Q$ silicon nitride (SiN) microcavity can be stabilized to within tens of MHz of any desired value and can be tuned repeatably by simply adjusting an electrical voltage.
+The frequency noise of such a microcavity is shown to be impervious to environmental drift and thermal crosstalk and limited only by $1/f$ flicker noise in the control electronics.
+We demonstrate that a one-time calibration of the resonance wavelength versus thermometer resistance is sufficient to repeatably tune the cavity to any desired resonance wavelength within the calibration range using thermometric control alone for at least a few days, removing the requirement for optical referencing or monitoring.
+By locking a DFB laser to the stabilized microresonator, we demonstrate that it can serve as a fully-integrated photonic frequency reference for telecommunications transceiver applications.
+The locked laser’s emission wavelength remains within $\pm 0.5$ pm of the mean over $50$ hours of measurement duration, outperforming many commercial DFB and wavelength-locker-based laser systems.
+Finally, we apply integrated thermometry to demonstrate a fully-thermally-stabilized Kerr comb that remains stably modelocked in the presence of strong thermal crosstalk.
+We envision this work to enable the realization of highly robust and field-deployable linear and nonlinear photonic devices for classical and quantum applications, and integrated Kerr combs that can potentially operate indefinitely.
+I Results
+Figure 1(a) conceptually illustrates integrated thermometry.
+A thin-film metallic resistor with a temperature-dependent resistance placed directly above a high-$Q$ microcavity acts as an on-chip resistance thermometer.
+Due to the low heat capacity of the thin-film resistor, small heat fluxes translate to large and observable changes in its temperature.
+The current-voltage-resistance (I-V-R) characteristics displayed in Fig.
+1(b) show a non-linear I-V relationship, and a quadratic dependence of R on V, indicating a linear dependence of R on the dissipated power.
+This behavior is similar to that of a filament lamp, which exhibits lower (higher) resistance at lower (higher) voltages due to lower (higher) filament temperature from Joule heating.
+Figure 1(c) shows that the thermometer resistance also depends quadratically on the voltage applied across the second heater, demonstrating that temperature changes induced by external heat sources can be measured by simply monitoring the resistance.
+Extended Data Fig.
+2 shows an independent measurement of the temperature-dependent resistance [52].
+We combine the temperature measurement scheme described above with a scanning-laser optical probe to perform rapid in-situ measurements of the thermo-optic coefficient of a silicon-nitride (SiN) microresonator’s waveguide mode.
+The free-spectral range (FSR) is 76 GHz, and the loaded $Q$ is approximately $3\times 10^{6}$.
+The measurement procedure is described in Methods.
+Figure 1(d) shows the measured resonance frequency shift and the corresponding change in thermometer resistance, illustrating a strong correlation.
+Figure 1(e) plots the resonance frequency shift versus the measured temperature change.
+The slope of the linear fit is used to extract the thermo-optic coefficient of the fundamental spatial mode of the SiN waveguide, $dn_{\mathrm{eff}}/dT=2.45\times 10^{-5}$ K-1.
+This value is approximately equal to the thermo-optic coefficient $dn/dT$ of SiN due to the high spatial confinement offered by the 2100-nm waveguide and is consistent with values measured using other techniques [53, 54, 55].
+Figure 2(a) illustrates the experimental schematic for investigating the long-term frequency stability of the microresonator using a piezo-controlled scanning probe laser.
+The absolute frequency calibration procedure is described in Methods.
+The microcavity resonance frequency is averaged over and recorded once every 10 seconds over the period of 24 hours.
+Three such measurements are performed, each under the following conditions, (i) ‘free-running’: the cavity is subject only to ambient temperature fluctuations, (ii) ‘open-loop’: additional perturbations are induced via a voltage-summing circuit connected to the heater to mimic thermal crosstalk from co-packaged electronics, and (iii) ‘stabilized’: the microresonator’s temperature is actively stabilized by applying proportional-integral-derivative (PID) feedback to the heater based on the voltage drop across the resistance thermometer connected to a low-noise current source.
+The strength of the perturbation applied is denoted by the empirical parameter $\xi$: for $\xi$=1, the root mean-squared (RMS) fluctuation of the resonance frequency equals the short-term resonance full-width-at-half-maximum (FWHM) linewidth, which is 75 MHz for the microcavity used in this work.
+A circuit schematic of the current source and its noise performance are provided in Supplementary Fig.
+S1.
+The thermal design of the dual-resistor scheme is explained in Extended Data Fig.
+1.
+Figure 2(b) displays the measured resonance frequency fluctuations.
+Without stabilization, ambient noise causes a significant resonance drift, even in a temperature-stabilized laboratory, as seen in the ‘free-running cavity’ trace.
+The resulting standard deviation of the resonance frequency is $\approx 341$ MHz.
+The additional perturbation applied ($\xi=7.55$) drives faster-timescale fluctuations seen in the ‘open-loop cavity’ trace.
+The measured open-loop standard deviation is $\approx 470$ MHz.
+Supplementary Figs.
+S2 and S3 provide the temperatures measured by the resistance thermometer for the free-running and open-loop cases, which show strong correlations with the resonance frequency shifts.
+Finally, the stabilized cavity exhibits a highly stable resonance frequency with significantly weaker fluctuations (standard deviation $\approx 13.7$ MHz, a $34\times$ reduction from the open-loop scenario) even in the presence of additional perturbation ($\xi=7.55$).
+Figure 2(d) visualizes the time-domain data as histograms (bin size =5=5 MHz).
+Notably, even in the free-running case, the resonance frequency histogram does not exhibit a well-defined peak.
+Figure 2(c) shows the Allan deviation (ADEV) versus averaging time computed from the time-domain data.
+The stabilized cavity offers a significantly lower ADEV compared to the free-running and open-loop cases.
+The flat shape of the stabilized cavity’s ADEV closely matches that of pure $1/f$ noise, indicating that unbounded random-walk resonance frequency drifts arising from environmental and crosstalk perturbations are effectively canceled out by the stabilization method, and that residual fluctuations are bounded and limited only by noise in the control electronics used in this work.
+Supplementary Fig.
+S4 provides ADEV comparison measurements demonstrating that the stabilized microresonator offers superior long-term frequency stability compared to many commercially available telecommunication lasers [56, 37] available in our laboratory, demonstrating its potential utility as a highly practical integrated reference cavity.
+The long-term repeatability of the microcavity’s tunable resonance wavelength is illustrated by the demonstration of a resonance-wavelength thermometric lookup table.
+A one-time calibration measurement of the absolute resonance wavelength versus the thermometer resistance is performed using the setup described in Fig.
+3a.
+The current source output is monitored by measuring the voltage across a series $50.0$ $\Omega$ bulk termination resistor using a precision 6.5-digit digital multimeter (DMM; Agilent 34401A).
+The thermometer voltage is measured by subtracting the voltage across the termination from the total voltage across the current source, which in stabilized operation equals the PID set-point voltage.
+The thermometer resistance calculated as the ratio of these two quantities also includes a small contribution from contact resistance of the spring-loaded electrical contact probes (PicoProbe 40A-GS-150-P) employed in our experimental setup.
+The results of the one-time calibration are shown in Fig.
+3b.
+The linear fit extracted from this calibration is used in subsequent measurements to compute the desired thermometer resistance to achieve any desired resonance wavelength within the calibration range.
+A computer-generated string of $10,000$ pseudorandom “target” resonance wavelengths is fed into a thermometric control loop implemented using a field-programmable gate array (FPGA), which calculates the desired thermometer resistance corresponding to a target resonance wavelength using the calibration fit, and stabilizes the microresonator to the computed temperature.
+The resonance wavelength is then measured and compared with the target value.
+Each such “measurement” takes approximately five seconds.
+Two separate sets of 10,000 measurements each are performed.
+A zoomed-in portion of the first set of 10,000 measurements, which concluded 12 hours after the calibration and is displayed in Fig.
+3c, shows close agreement between the measured and target resonance wavelengths.
+Figure 3(d) shows histograms of the absolute-wavelength error (bin size =0.1=0.1 pm), defined as the target minus measured wavelengths.
+For the first set, a root-mean squared error (RMSE) is $0.11$ pm is observed.
+A systematic error, resulting in a larger RMSE of $0.77$ pm, was observed in the second set carried out 3 days after the calibration.
+It is hypothesized to be stemming from drift in the contact resistance of the spring-loaded probes.
+The repeatability of the thermometrically controlled microcavity’s resonance wavelength is further demonstrated by programming the resonance wavelength to trace a predetermined pattern.
+When the heater is actuated without feedback control, in the presence ($\xi=3.77$) as well as absence ($\xi=0$) of additional perturbation, significant long-term drift is observed over the $10$-hour measurement duration (Figs.
+4(a) and (b)).
+On the other hand, when the microresonator is stabilized and the set-point voltage is tuned instead, long-term drift is canceled out, and the resonance wavelength exhibits a highly repeatable relationship with the set-point voltage, as shown in Fig.
+4(c).
+The frequency stability of the microresonator is transferred to a fiber-coupled DFB laser using the Pound-Drever-Hall (PDH) method.
+For an integrated laser scheme, self-injection locking may be more suitable [22].
+The DFB emission wavelength is recorded once every 10 seconds using a wavelength meter.
+As shown in Fig.
+5(a), the free-running DFB experiences significant wavelength drift due to ambient temperature fluctuations, with a standard deviation of $11.5$ pm (i.e., $1.43$ GHz) over $42$ hours of measurement time.
+In comparison, the locked DFB is highly stable, with a peak-to-peak variation in emission wavelength that $\pm 0.5$ pm of the mean for over $50$ hours, and a standard deviation of $0.24$ pm (i.e., $30$ MHz).
+The stability enhancement factor is $\approx 48\times$.
+The long-term stability reported here is superior to many commercially available DFB and wavelength-locker-based laser systems.
+Figure 5(b) shows that the ADEV of the locked DFB is significantly lower and is comparable to the ADEV of the stabilized microresonator displayed in Fig.
+2.
+Due to the large integral gain of the PDH servo, a lower ADEV is observed at smaller averaging times.
+The higher ADEV at larger averaging times is attributed to insufficient gain of the PDH servo at very low frequencies.
+Finally, we show in Fig.
+6(a) an approach to stabilize a modelocked Kerr comb, which is typically highly sensitive to thermal perturbations.
+A $1550$-nm pump laser is amplified before coupling into the photonic chip (pump power =150=150 mW).
+The thermometer current source is modulated to generate a single-soliton Kerr comb [49].
+The comb power is measured using a photodiode after filtering the pump line using a fiber-optic notch filter.
+The stability of the Kerr comb is studied by applying a white-noise perturbation (bandwidth =20=20 Hz) via the outer on-chip heater and measuring the on-chip temperature and the output comb power.
+The repetition-rate noise is measured by generating a microwave using a $100$-GHz bandwidth photodetector and characterizing it using a phase noise analyzer (PNA; Rohde & Schwarz ®  FSWP50) via an 8:1 harmonic mixer.
+For $\xi=1.51$ and without feedback correction, the modelocked Kerr comb transitions to a chaotic comb state almost immediately, as seen in Fig.
+6(c).
+The thermometer voltage plot shows a sudden increase, indicating an abrupt jump to a state with higher intracavity power, such as a chaotic comb.
+This is corroborated by the random fluctuations exhibited by the comb power and the thermometer voltage following the jump.
+The smaller amplitude of fluctuations observed by the thermometer following the jump result from an increase (decrease) in temperature at the thermometer due the perturbation and its consequent secondary effect, namely decrease (increase) in thermometer temperature due to decrease (increase) in intracavity optical power and self-heating.
+This is caused by the effectively blue-detuned pump of the chaotic comb and is not applicable to the modelocked comb due to its effectively red-detuned pump.
+In the stabilized case ($\xi=7.55$; Fig.
+6(d)), the PID correction voltage cancels the applied perturbation, leading to a stable temperature and comb power.
+Whereas the Kerr comb is pushed out of its existence range immediately in the open-loop ($\xi=1.51$) scenario, closed-loop operation was observed to support modelocking for over 20 minutes, limited only by the drift in the input fiber-to-chip optical coupling.
+In the absence of coupling drift, such as in packaged photonic devices, cold-cavity detuning can be made impervious to external thermal fluctuations.
+Thermometric control can then enable Kerr-comb-based photonic devices that potentially remain modelocked indefinitely.
+Lastly, Fig.
+6(e) shows the measured microwave single-sideband (SSB) frequency noise spectra.
+For the ‘open-loop’ scenario, the noise strength is reduced ($\xi=0.75$) to maintain modelocking.
+The microwave noise for the stabilized case is >20>20 dB below the open-loop case despite the stronger added-noise strength ($\xi=7.55$).
+At offset frequencies below $\sim 1$ Hz, the stabilized cavity’s microwave exhibits lower noise than the free-running microwave since ambient drift is more prominent on longer timescales.
+A comparison of the measured soliton repetition rates is presented in Supplementary Fig.
+S5.
+II Discussion
+In summary, we have demonstrated a powerful new approach based on fully integrated thermometry to stabilize the resonance frequency of high-$Q$ microresonators without requiring any photodetection.
+Thermometric stabilization mitigates environmental as well as crosstalk-induced thermal perturbations, delivering a highly stable microresonator whose residual frequency fluctuations are only limited by $1/f$ noise in the control electronics.
+Realizing temperature-stabilized microresonators that are TRN-limited on long timescales is an exciting topic for future research, as it offers the prospect of precision metrology on a chip.
+We demonstrate that after a one-time calibration, the absolute resonance wavelength can be tuned to any desired value within the calibration range via thermometric control, with measured absolute-wavelength RMSEs $0.11$ pm and <0.77<0.77 pm after $12$ hours and $3$ days following the calibration, respectively.
+The emission wavelength of a DFB laser frequency-locked to the microresonator remains within $\pm 0.5$ pm of the mean over $50$ continuous hours, which is a $48\times$ improvement compared to free-running operation.
+This performance is superior to many commercial DFB and wavelength-locker-based laser systems.
+Finally, we stabilize a Kerr comb against strong thermal perturbations, offering a path toward chip-based comb devices that can potentially remain modelocked indefinitely in noisy environments.
+Although our demonstration was focused on SiN microresonators, our technique is readily applicable to other dielectric and semiconductor photonic materials.
+Removing the requirement of drop-ports and photodetection for stabilization can significantly reduce design complexity and fabrication costs of practical photonic devices.
+Finally, the concepts described here can be applied to several other photonic integrated systems, including microresonator arrays and Mach-Zehnder interferometer (MZI) meshes for topological photonics [57] and photonic computing [58, 59].
+III Acknowledgments
+The authors acknowledge support from the Army Research Office under Cooperative Agreement Number W911NF-24-2-0204 (M.L.
+and A.L.G.), the Air Force Office of Scientific Research under FA9550-20-1-0297 (M.L.
+and A.L.G.), and the Defense Advanced Research Projects Agency under HR0011-19-2-0014 (M.L.
+and A.L.G.).
+The authors wish to acknowledge Kaylx Jang and Michael Cullen from the group of Prof.
+Keren Bergman for providing the photograph of a co-packaged silicon photonics device, previously reported in Ref.
+[60], included in Fig.
+1.
+The authors thank Vivian Zhou from the group of Prof.
+Michal Lipson for assistance in setting up the vacuum system used in the temperature coefficient of resistance measurements.
+IV Author Contributions
+S.K.D.
+and A.L.G.
+conceived the project.
+S.K.D.
+and Y.Z.
+performed the experiments.
+S.K.D.
+and G.R.B.
+performed the design and characterization of the resistance thermometer.
+S.K.D., Y.Z., and A.L.G.
+performed the data analysis with input from all authors.
+K.J.M.
+fabricated the devices under the supervision of M.L.
+S.K.D.
+and A.L.G.
+wrote the manuscript with feedback from all authors.
+M.L.
+and A.L.G.
+supervised the project.
+V Competing Interests
+S.K.D., Y.Z., K.J.M., G.R.B., M.L., and A.L.G.
+are inventors on a U.S.
+provisional patent application filed by The Trustees of Columbia University in the City of New York (application number 63/830,566) on 26 June 2025 that covers ultrastable microresonators using integrated thin-film resistance thermometers and thermistors.
+Figure 1: Integrated thermometry.
+a, Conceptual illustration of integrated thermometry applied to stabilize a high-Q monolithic microresonator against thermal fluctuations arising from ambient heat sources and crosstalk from other thermally-tuned devices on the same chip.
 A thin-film metallic (Platinum) resistor, typically used as a microheater, exhibits a temperature-dependent resistance due to the metal’s natural temperature-dependent resistivity.
@@
 A second identical resistor is used as a heater to perform active stabilization purely using the resistance thermometer, thereby eliminating the need for optical probing for stabilization.
-Image for data center sourced from Ref.
+Image for data center sourced from Adobe Stock, Ref.
 [61].
+Image courtesy for picture of co-packaged device: Kaylx Jang.
 b, Current-voltage-resistance (I-V-R) characteristics of the resistance thermometer fabricated in this work.
@@
 c, The measured thermometer resistance (lavender dots) also exhibits a quadratic dependence (green line) on the voltage applied across the (second) heater, indicating a linear dependence on power dissipated across the heater.
-d, Resonance frequency shift (green circles) of a high-Q microcavity of free-spectral range (FSR) 76 GHz, measured using a piezo-tuned tunable laser source, and measured change in thermometer resistance (gold line), for a sinusoidal perturbation applied via the heater.
+d, Resonance frequency shift (green circles) of a high-Q microcavity of free-spectral range (FSR) 76 GHz, measured using a calibrated piezo-tuned probe laser, and measured change in thermometer resistance (gold line), for a sinusoidal perturbation applied via the heater.
 The cavity temperature and resonance frequency exhibit a strong anti-correlation.
@@
 The linear fit provides an in-situ measurement of the thermo-optic coefficient of the fundamental spatial mode of the cavity waveguide.
-The temperature measurement scheme can be utilized together with optical measurements to perform rapid in-situ measurements of the thermo-optic coefficient of the microcavity waveguide’s spatial mode at any desired wavelength.
-In order to demonstrate this, a small-amplitude 2-Hz sinusoidal voltage waveform is applied to the heater, while monitoring the electrical resistance of the thermometer.
-The free-spectral-range (FSR) of the SiN microresonator is 76 GHz, its loaded $Q$ is approximately $3\times 10^{6}$, and the resonances in the telecommunications C-band are critically coupled.
-The width of the waveguide is 2100 nm.
-A tunable laser source centered at 1560 nm is swept at 200 Hz, using its internal piezo-electric module, across a single microcavity resonance.
-By synchronously monitoring the transmitted optical output using a photodiode connected to a fast oscilloscope, the resonance frequency shift of the cavity is measured.
-Simultaneously, the on-chip temperature change is also monitored on the same oscilloscope by sending a constant low-noise DC current through the resistance thermometer and measuring the change in voltage across it.
-Figure 1(d) shows the measured resonance frequency shift (green circles; left y-axis) and the corresponding change in thermometer resistance (golden solid line; right y-axis), indicating a clear correlation between the two.
-The measured change in thermometer resistance is converted to a change in temperature using the following equation,
-$R(T)=R_{0}(1+A\Delta T)$ (1)
-where $R_{0}$ is the resistance at room temperature, $\Delta T$ is the change in temperature, and $A$ is the temperature coefficient of resistance (TCR).
-The coefficient $A$ was measured independently to be $A=0.00172$.
-The measurement procedure is detailed in Methods.
-Figure 1(e) shows the data presented in Fig.
-1(d), now with the resonance frequency shift plotted against the measured change in temperature.
-The slope of the linear fit is then used to extract the thermo-optic coefficient of the fundamental spatial mode of the waveguide, $dn_{eff}/dT=2.45\times 10^{-5}$ K-1.
-This value is also approximately equal to the thermo-optic coefficient $dn/dT$ of SiN due to the high spatial confinement offered by the fundamental mode of the 2100-nm waveguide, and is consistent with values published previously that employ different measurement techniques [46, 47, 48].
-III Fully-Integrated Photonic Frequency Reference
-The frequency stability of the microresonator is investigated, as depicted in Fig.
-2(a), by piezo-scanning a probe laser across a cavity resonance at 200 Hz, and recording the resonance frequency position once every ten seconds.
-Absolute wavelength calibration is performed by beating the tunable laser with a low-drift fixed-wavelength reference laser on a fast photodiode, and simultaneously calibrating the reference laser’s wavelength using a benchtop wavemeter (Bristol Instruments 238A).
-This yields a calibrated absolute cavity resonance frequency, which is recorded continually once every 10 seconds over the course of 24 hours.
-The microcavity’s frequency stability is measured under the following three conditions: (i) ‘free-running’, in which the cavity is subject only to ambient fluctuations (such as change in laboratory temperature), (ii) ‘open-loop’, in which an additional perturbation is induced via a voltage summing circuit connected to the heater, in order to emulate real-world use, and (iii) ‘stabilized’, in which the microresonator’s temperature is actively stabilized.
-The strength of the perturbation added is denoted by the parameter $\xi$, which is defined as follows.
-For $\xi$=1, the root mean-squared (RMS) fluctuation of the cavity resonance frequency equals its short-term resonance full-width-at-half-maximum (FWHM) linewidth, which is 75 MHz for the microcavity used in this work.
-The cavity is stabilized by sensing the temperature in real-time by sending a low-noise electrical current through the resistance thermometer and measuring voltage changes across it, and by performing active feedback to the heater using a low-noise proportional-integral-derivative (PID) circuit.
-A schematic of the current source fabricated for this work and its noise performance are provided in Supplementary Information.
-The thermal design of the dual-resistor scheme described here is explained further in Methods.
 Figure 2: Long-term stabilization of the absolute resonance frequency of a high-Q SiN microresonator using integrated thermometry.
 a, Simplified experimental schematic, depicting a tunable probe laser that is piezo-scanned across a microcavity resonance.
-Absolute frequency calibration of the scan is performed by producing a heterodyne note between the tunable laser and a stable reference laser, and simultaneously tracking the reference laser’s wavelength drift using a highly-stable benchtop wavelength meter.
-The heterodyne beat note as well as the cavity resonance line-shape are measured synchronously with the piezo sweep using an real-time oscilloscope connected to a computer interface for periodic data acquisition every 10 seconds.
-For $\xi$=1, the pseudorandom perturbation applied via the outer heater yields a root mean-squared (RMS) of the induced cavity resonance frequency fluctuation equals its short-term resonance full-width-at-half-maximum (FWHM) linewidth, which is 75 MHz for the microcavity fabricated in this work.
+Absolute frequency calibration of the scan is performed by producing a heterodyne note between the tunable laser and a stable reference laser, and simultaneously tracking the reference laser’s wavelength drift using a benchtop wavelength meter.
+The heterodyne beat note as well as the cavity resonance line-shape are measured synchronously with the piezo sweep using a real-time oscilloscope connected to a computer interface for periodic data acquisition every 10 seconds.
+For $\xi$=1, the pseudorandom perturbation applied via the heater yields a root mean-squared (RMS) of the induced cavity resonance frequency fluctuation that is equal to the cavity resonance full-width-at-half-maximum (FWHM) linewidth, which is 75 MHz for the microcavity fabricated in this work.
 b, Drift in the microcavity’s calibrated resonance frequency over the duration of 24 hours, for free-running ($\xi$=0), open-loop ($\xi$=7.55), and stabilized ($\xi$=7.55) operation.
@@
 d, Histogram plots depicting the fraction of time over 24 hours (y-axis) that the microcavity’s absolute resonance frequency drift lies within a given frequency bin with a width of 5 MHz.
-Figure 2(b) displays the measured resonance frequency drift of the microresonator for the three cases described above over a period of 24 hours.
-In the absence of any stabilization, ambient noise alone causes a significant resonance drift, even in temperature-stabilized laboratory environments.
-This is shown by the slow cyclical pattern in the ‘free-running cavity’ data points, with an standard deviation of resonance frequency variation of $\approx 341$ MHz.
-Additional perturbation, such as from thermal crosstalk in practical applications, adds to the drift in the form of faster-timescale fluctuations, as evidenced by the ‘open-loop cavity’ data points, exhibiting a standard deviation of resonance frequency variation of $\approx 470$ MHz.
-The Supplementary Information provides the corresponding real-time temperatures measured by the resistance thermometer for the free-running and open-loop cases, which shows excellent correlation with the resonance frequency shift.
-Finally, the stabilized cavity exhibits a highly stable resonance frequency with a considerably lower drift even in the presence of significant added perturbation ($\xi=7.55$).
-The standard deviation of the measured resonance frequency variation of the stabilized microresonator is $\approx 13.7$ MHz, corresponding to a 34×34$\times$34 × reduction compared to the open-loop scenario.
-The time-domain measurement data is better visualized as histograms, shown in Fig.
-2(d).
-The bin-size along the x-axis is 5 MHz.
-Remarkably, even in the absence of additional perturbations, i.e., in the free-running case, the cavity resonance frequency histogram does not consist of a single, well-defined narrow peak.
-Figure 2(c) shows the numerically-computed Allan deviation (ADEV).
-As expected, the stabilized cavity offers a significantly lower ADEV compared to the free-running and open-loop cavities.
-The shape of the ADEV of the stabilized cavity closely matches that of pure $1/f$ noise, indicating that unbounded random walk resonance frequency drifts arising from environmental and crosstalk perturbations are effectively canceled out by the stabilization method employed and that the residual fluctuations are bounded, limited only by noise in the control electronics used in this demonstration.
-This is evidenced by the flat shape of the ADEV versus averaging time, compared to the free-running and open-loop cavities, which exhibit non-stationary unbounded random walks, shown by the increasing ADEV at large averaging times.
-The stabilized microresonator also offers significantly higher long-term stability compared to temperature-stabilized DFB lasers, which are commonly used in chip-based telecommunications applications [49, 30], making it a highly practical fully-integrated frequency reference cavity.
-Finally, long-term repeatability of the microcavity’s tunable absolute resonance wavelength is assessed as follows.
-To assess the long-term repeatability of a thermally-tuned cavity only subject to ambient thermal noise (i.e., $\xi=0$), the voltage applied to the outer heater is swept in discrete steps so as to trace the logo of Columbia University, and the microcavity’s absolute resonance wavelength is synchronously measured using the scanning probe laser setup described in Fig.
-2(a).
-This measurement is repeated in the presence of additional perturbations to mimic external thermal fluctuations in practical photonic devices ($\xi=3.77$).
-No feedback is applied in the first two configurations.
-Finally, the feedback loop is closed, and thermal tuning is performed by simply changing the set-point of the PID loop.
-The results are shown in Fig.
-3.
-Figure 3: Demonstration of tunability and long-term repeatability of the microresonator’s absolute resonance wavelength.
+Figure 3: Demonstration of thermometric absolute-resonance-wavelength lookup table.
+a., Simplified experimental schematic, depicting a tunable probe laser that is piezo-scanned across a microcavity resonance.
+The absolute resonance wavelength is measured by producing a heterodyne beat between the tunable laser and a stable reference laser, and simultaneously tracking the reference laser’s wavelength using a highly-stable benchtop wavelength meter.
+The current produced by the low-noise current source is monitored by measuring the voltage across a series $50.0$ $\Omega$ bulk resistor using a precision 6.5-digit digital multimeter (DMM).
+The PID loop is implemented using a commercial FPGA, and the set-point voltage is input via a computer interface.
+b.
+Calibration measurement of the absolute resonance wavelength vs the thermometer resistance.
+The linear fit extracted from this one-time calibration is utilized in subsequent measurements to compute the desired thermometer resistance to achieve any desired resonance wavelength within the calibration range.
+c.
+Demonstration of arbitrary absolute-wavelength tuning of the microresonator via thermometry.
+A string of 10,000 pseudorandom “target” wavelengths (red hollow circles) are provided as inputs to the thermometric control system illustrated in a., which computes the desired temperature for each target wavelength and stabilizes the microresonator to the computed temperature.
+The resonance wavelength is measured (blue-magenta filled circles) and compared with the target wavelengths.
+A zoomed-in portion of 100 consecutive measurements out of 10,000 is shown.
+d.
+Histograms showing the deviation between the target and measured absolute resonance wavelengths following a one-time calibration.
+The histogram bin size is 0.1 pm.
+The first set of 10,000 measurements, which concluded 12 hours after the calibration run, exhibits a root-mean squared absolute-wavelength error of 0.11 pm.
+The second set of measurements three days later shows a similar narrow distribution of absolute-wavelength error, but exhibits a systematic error, resulting in a 0.77 pm.
+Figure 4: Tunability and long-term repeatability of the microresonator’s absolute resonance wavelength.
 The resonance frequency is programmed to trace a pre-defined shape by blind tuning of the power supplied to the outer heater in the presence of ambient noise alone ($\xi$=0; shown in a), in the presence of additional perturbation ($\xi$=3.77; shown in b), and by tuning of the set-point for temperature stabilization in the presence of ambient noise as well as additional perturbation ($\xi$=3.77; shown in c).
@@
 The stabilized tuning, on the other hand, yields a highly repeatable performance even in the presence of strong ambient and external noise.
-When tuned with the heater alone, i.e., the most common way of tuning microresonators, applying the same voltage does not lead to the same resonance wavelength over the measurement duration of 10 hours, as evidenced by the drift in the patterns in Figs.
-3(a) and (b).
-Similar to the results presented in Fig.
-2, additional perturbations lead to “short-term” fluctuations superimposed on top of the long-term drift caused by environmental drift.
-However, in the stabilized case in which the set-point of the feedback loop is tuned, active feedback based on the resistance thermometer’s “reading” cancels out ambient drift as well as the perturbation-induced fluctuations, and the microcavity resonance wavelength exhibits significantly better repeatability.
-The long-term frequency stability of the thermometrically stabilized microresonator is demonstrated further by stabilizing the frequency of a laser.
-A DFB laser is locked to a microcavity resonance using the Pound-Drever-Hall (PDH) method, and its emission center wavelength is recorded every 10 seconds using a bench-top wavelength meter.
-We employ the PDH method in this experiment due to its insensitivity to drift in the input fiber-to-chip optical coupling.
-In a fully-integrated system, however, the frequency stability of the microresonator is easily transferred to an integrated laser via mechanisms such as self-injection locking [50].
-The results of the frequency stabilization demonstration are shown in Fig.
-4.
-Figure 4(a) shows the measured drift in laser center wavelength for the free-running and PDH-locked DFB lasers, in blue and maroon respectively.
-The mean emission wavelength for the two cases are approximately 1551.2 nm.
-The free-running DFB experiences significant wavelength drift arising from ambient temperature fluctuations alone, amounting to a standard deviation of $11.5$ pm (i.e., $1.43$ GHz) over the span of $42$ hours of measurement time.
-In comparison, the locked DFB emits at a highly stable center wavelength, with a peak-to-peak variation that lies within ±$\pm 0.5$ pm of the mean over $50$ hours of measurement time, and exhibits a standard deviation of $0.24$ pm.
-The stability enhancement factor is 48$\approx$ 48$\times$ 48 ×.
-The long-term stability measured here is significantly superior to many commercially-available DFB laser system, and makes the scheme presented here ideal for the realization of practical photonic transceivers for dense wavelength division multiplexing (DWDM) applications.
-Figure 4(b) shows that the computed ADEV of the locked DFB is significantly lower ADEV than the free-running DFB, and is comparable to the ADEV of the stabilized microresonator displayed in Fig.
-2.
-By using a large integral gain in the feedback servo of the PDH, a slightly lower ADEV is observed at smaller averaging times.
-The slightly higher ADEV at larger averaging times is attributed to insufficient gain of the PDH feedback servo at very low frequencies.
 Figure 5: Demonstration of long-term laser frequency stabilization to the microresonator.
@@
 a, Time-domain comparison of center wavelength drift of free-running and PDH-locked configurations of a DFB laser.
-The locked laser’s center wavelength remains within a ±$\pm 0.5$ pm window of the mean, indicated by the dashed lines, for 50 hours.
+The locked laser’s center wavelength remains within a $\pm 0.5$ pm window of the mean, indicated by the dashed lines, for 50 hours.
 The free-running laser is only subject to ambient temperature drift in a temperature-stabilized laboratory environment.
@@
 The slightly higher ADEV at larger averaging times is attributed to insufficient gain of the PDH feedback servo at very low frequencies.
-IV Fully-thermally-stabilized Kerr Comb
-Finally, we show that the fully-integrated thermometry approach described in this work can stabilize a modelocked Kerr comb, which is typically highly sensitive to perturbations, for long periods.
-The experimental schematic is shown in Fig.
-5(a).
-A pump laser at approximately 1550 nm is amplified before coupling into the microcavity device.
-The off-resonance optical power output is 150-mW.
-The low-noise current source connected to the resistance thermometer is modulated with a burst waveform to generate a single-soliton Kerr comb, shown in blue in the optical spectrum analyzer (OSA) measurements in Fig.
-5(b).
-A portion of the out-coupled Kerr comb is sent through a fiber-optic notch filter centered at the pump wavelength, and the transmitted output, shown in red in Figure 5(b), is used to track the comb power using a photodiode.
-Phase-noise characterization is performed by generating a microwave signal at the repetition rate of the comb (76 GHz) via a high-speed photodetector (bandwidth 100 GHz), and measuring the microwave phase noise using a phase noise analyzer (PNA; Rohde & Schwarz ®  FSWP50).
-Since the microwave frequency (76 GHz) is beyond the specified bandwidth (50 GHz) of the PNA used in this work, an 8:1 harmonic mixer (Rohde & Schwarz ®  RPG FS-Z110) is employed to down-convert the microwave frequency to within the PNA’s bandwidth.
-The stability of the Kerr comb is studied by applying a perturbation (white-noise voltage, bandwidth $=20$ Hz) via the outer on-chip heater and measuring the real-time on-chip temperature and the power and phase noise of the comb.
-For a noise strength of $\xi=1.51$, without any feedback correction, the modelocked Kerr comb transitions to a chaotic comb state almost immediately, as shown in Fig.
-5(c), where the top plot shows the applied thermal perturbation voltage signal in the time-domain.
-The thermometer voltage plot, for a constant low-noise current sent through the resistance thermometer, shows a sudden increase in the cavity temperature, indicating a sudden jump to a state with higher intracavity power.
-This is explained by the bottom plot, which shows that the comb power (which is measured by filtering out the pump line) is no longer stable and exhibits random fluctuations driven by the perturbation voltage.
-These fluctuations are also seen in the thermometer voltage following the jump.
-The smaller amplitude of fluctuations observed by the thermometer is a result of an increase (decrease) in temperature at the thermometer due the perturbation and the consequent secondary effect, namely decrease (increase) in temperature at the thermometer due to decrease (increase) in intracavity optical power and self-heating.
-This is a result of the inverse relationship between cavity temperature and intracavity power for the effectively blue-detuned pump of the chaotic comb state, which does not apply for the modelocked comb state, since the pump is effectively red-detuned, and the resistance thermometer therefore remains sensitive to external perturbations, making it an efficient tool in stabilizing modelocked comb states.
 Figure 6: Fully-thermally stabilized Kerr optical frequency comb.
@@
 The microwave generated in the stabilized case has considerably lower noise than the open-loop case despite the much stronger noise.
-The stability of the closed-loop system is studied by repeating the same measurement sequence as above.
-However, in this case, the noise strength is increased five-fold to $\xi=7.55$.
-As seen in Fig.
-5(d), the PID correction voltage cancels out the applied perturbation, leading to a stable temperature and Kerr comb power even in the presence of strong perturbations.
-Whereas the Kerr comb is pushed out of its existence range immediately in the open-loop ($\xi=1.51$) scenario, closed-loop operation was observed to support modelocking for over 20 minutes, limited only by the drift in the input fiber-to-chip optical coupling, paving the way for Kerr-comb-based photonic devices that can operate in the desired modelocked state indefinitely.
-Lastly, Fig.
-5(e) shows the comparison of the microwave single-sideband (SSB) frequency noise.
-For the ‘open-loop’ scenario, the noise strength is reduced to $\xi=0.75$ to keep the comb in the modelocked regime of operation.
-For the ‘stabilized’ scenario, the noise strength is maintained at $\xi=7.55$.
-As the yellow and purple solid curves show, the frequency noise of the microwave generated by the stabilized cavity is over 20 dB below that of the microwave generated by the open-loop cavity.
-At lower offset frequencies (i.e., over longer timescales), the stabilized cavity’s microwave exhibits a lower noise than the free-running cavity’s microwave since ambient drift is more prominent on longer timescales.
-V Conclusion
-In conclusion, we demonstrate a powerful new approach based on fully-integrated thermometry to stabilize the absolute resonance frequency of high-$Q$ silicon nitride microcavities without requiring more complicated monitoring systems such as photodetection.
-The thermometric stabilization mitigates environmental drift as well as externally-induced perturbations effectively, and the Allan deviation of the stabilized microcavity’s resonance frequency remains nearly flat versus averaging time, in contrast to the unbounded random walks exhibited by the free-running and open-loop cases, evidenced by the increasing Allan deviation at large averaging times.
-By locking a DFB laser to the stabilized microresonator, we demonstrate that the center-wavelength stability of the laser is <±$<\pm 0.5$ pm from the mean over the measurement duration of $50$ continuous hours, which is a 48×48$\times$48 × improvement compared to the free-running laser.
-This work paves the way for highly compact, fully-integrated and field-deployable frequency references for telecommunications applications.
-When pumped with a high-power laser source to generate a Kerr comb, the thermometry-based stabilization technique described in this work makes the comb significantly more stable even in the presence of strong perturbations.
-The fully-thermally-stabilized comb demonstrated here offers a path toward comb-based high-$Q$ photonic devices that can operate indefinitely even in noisy environments.
-Although the demonstration in this work was limited to the SiN platform, this technique is readily generalizable to other photonic materials.
-Notably, previously-demonstrated on-chip temperature-sensing techniques, such as ones leveraging the photoconductive effect, require the presence of significant optical power in the cavity, and are limited to semiconductor platforms.
-The resistance-thermometer-based scheme here poses no such requirement.
-Finally, removing the requirement of additional drop-ports and photodetection for stabilization can reduce the design complexity and fabrication costs of practical photonic devices quite significantly.
-VI Acknowledgments
-The authors wish to acknowledge Kaylx Jang and Michael Cullen from the group of Prof.
-Keren Bergman for providing the photograph of a co-packaged silicon photonics device, previously reported in Ref.
-[51], included in Fig.
-1.
-The authors thank Vivian Zhou from the group of Prof.
-Michal Lipson for assistance in setting up the vacuum system used in the temperature coefficient of resistance measurements.
-VII Data Availability
-The data that support the findings of this study are available from the corresponding author upon reasonable request.
-VIII Funding
-The authors acknowledge support from the Army Research Office under Cooperative Agreement Number W911NF-24-2-0204, the Air Force Office of Scientific Research (FA9550-20-1-0297), and the Defense Advanced Research Projects Agency (HR0011-19-2-0014).
-Ultrastable nanophotonic microcavities via integrated thermometry: Methods
-I Thermal Design for Integrated Thermometry
-Figure M1: Thermal design for integrated thermometry.
-a, Numerically simulated steady-state temperature profile for the scheme described in this work.
-The heater power is set to $100$ mW for the purposes of illustration.
-The initial temperature of the chip is $293$ K.
-b, Heater-thermometer spacing design trade-off, showing the core-thermometer temperature differential (left y-axis, in blue), and the heater power required for core heating of $0.1$ K (right y-axis, in dark red), versus heater-thermometer spacing.
-Proper thermal design is essential for realizing an on-chip temperature-stabilization scheme that is effective at stabilizing the resonance frequency of the high-$Q$ microresonator.
-Specifically, the heater must be placed at a radial separation from the microresonator such that the any thermal correction it applies leads to equal temperature changes at the resistance thermometer as well as the microresonator core.
-The right heater-thermometer spacing is determined prior to device fabrication by solving the following heat transfer equation in steady-state using the COMSOL Multiphysics® software:
-$\rho C_{p}\frac{\partial T}{\partial t}+\nabla\cdot\bm{q}=Q+Q_{ted}$ (M1)
-where $\rho$ is the material density, $C_{p}$ is the specific heat capacity at constant stress, $T$ represents the temperature, $q=-k\nabla T$ is the heat flux by conduction, $k$ is the thermal conductivity, $Q$ contains additional heat sources, and $Q_{ted}$ accounts for thermo-elastic damping in the material.
-All material thermal properties were set by the standard COMSOL material database for the cross-section of our devices.
-Figure M1(a) displays the cross-section of our device, and the steady-state spatial temperature profile of the stationary thermal simulation.
-The heat power is set to $100$ mW for the purposes of illustration.
-The initial temperature of the entire chip, and the boundary condition temperature for the bottom face of the chip, are set to $293$ K.
-Here, we denote the resulting steady-state temperature change at the center of the SiN core and at the center of the resistance thermometer as $\delta$Tcore and $\delta$Ttherm., respectively.
-Figure M1(b) shows the temperature differential $\delta$Tcore - $\delta$Ttherm.
-(left y-axis, in blue) as well as the heater power required to heat up the core (right y-axis, in dark red), which is a measure of heater efficiency, versus the heater-thermometer spacing.
-For small spacing, the heater-thermometer and heater-core separations are considerably different, leading to a large temperature differential.
-However, the close proximity allows for greater heater efficiency.
-For large spacing, the temperature differential drops, as the two separations are almost equal, while the heater efficiency worsens.
-The design employed in this work utilizes a $20$ $\mu$m heater-thermometer spacing, which leads to a very low temperature differential of $<5$ mK, making it effective in stabilizing the resonance frequency of the microcavity.
-The design principles described here can be readily applied to other integrated material platforms.
-II Device Fabrication
+Frequency-stable nanophotonic microcavities via integrated thermometry: METHODS
+I Device Fabrication
 The on-chip resistance thermometer demonstrated in this work is similar to devices that have been extensively used as resistive microheaters [48, 49].
@@
 The use of sputtered thin-films in our fabrication eliminates the need for processes requiring high-temperature and/or corrosive processing (e.g., electroplating).
-III Measurement of Temperature Coefficient of Resistance (TCR)
-Figure M2: Measurement of temperature coefficient of resistance of the integrated Platinum thermometer.
+II Calibrated Measurement of Absolute Resonance Frequency
+[…]
+III Thermal Design for Integrated Thermometry
+Proper thermal design is essential for realizing an on-chip temperature-stabilization scheme that is effective at stabilizing the resonance frequency of the high-$Q$ microresonator.
+Specifically, the heater must be placed at a radial separation from the microresonator such that the any thermal correction it applies leads to equal temperature changes at the resistance thermometer as well as the microresonator core.
+The right heater-thermometer spacing is determined prior to device fabrication by solving the following heat transfer equation in steady-state using the COMSOL Multiphysics® software:
+$\rho C_{p}\frac{\partial T}{\partial t}+\nabla\cdot\bm{q}=Q+Q_{ted}$ (E1)
+[…]
+[…]
+IV Measurement of Temperature Coefficient of Resistance (TCR)
+The integrated Platinum resistance thermometer, which allows the real-time measurement of the microresonator as described in the Main Text, is calibrated as shown in Fig.
+2(a).
+A test device is placed on a temperature-controlled chuck using a thermally-conductive adhesive in order to ensure proper thermal contact.
+The setup is placed and operated inside a vacuum probe station in order to minimize unwanted convective cooling.
+Two probes, connected to a Keithley $2400$ series SourceMeter®, are brought in electrical contact with the two ends of the Pt resistor under test (labeled as device under test (DUT) in Fig.
+2).
+At a given temperature of the chuck, the resistance $R$ of the DUT is measured by performing a current-voltage characterization.
+The voltage drop across the two contact pads is measured at various source currents, and a linear fit of the resulting data is obtained to extract $R$.
+This method is further explained in a prior manuscript [52].
+The source current values used are low enough (<1<1 mA) to avoid any self-heating of the resistor.
+This process is repeated for a range of chuck temperatures $T$.
+Figure 2(b) displays the results.
+Over the range of temperatures probed on this experiment, the resistance exhibits a linear trend.
+The temperature-coefficient of resistance is extracted by fitting the experimental data to Eq.
+E2, and is determined to be $A=0.00172$.
+We note that DUT described here has a shorter fabricated length compared to the resistance thermometer described in the Main Text, leading to the lower base resistance.
+V Measurement of Waveguide Thermo-Optic Coefficient
+A small-amplitude 2-Hz sinusoidal voltage is applied to the heater while monitoring the electrical resistance of the thermometer.
+Simultaneously, the shift in resonance frequency is probed using a scanning laser centered at 1560 nm, as described above.
+The free-spectral-range (FSR) of the SiN microresonator is 76 GHz, its loaded $Q$ is approximately $3\times 10^{6}$, and the resonances in the telecommunications C-band are critically coupled.
+The width of the waveguide is 2100 nm.
+The measured change in thermometer resistance is converted to a change in temperature using the following equation,
+$R(T)=R_{0}(1+A\Delta T)$ (E2)
+where $R_{0}$ is the resistance at room temperature, $\Delta T$ is the change in temperature, and $A$ is the temperature coefficient of resistance (TCR).
+The coefficient $A$ is measured independently to be $A=0.00172$.
+The measurement procedure is detailed in Extended Data Fig.
+2.
+Extended Data Figure 1: Thermal design for integrated thermometry.
+a, Numerically simulated steady-state temperature profile for the scheme described in this work.
+The heater power is set to $100$ mW for the purposes of illustration.
+The initial temperature of the chip is $293$ K.
+b, Heater-thermometer spacing design trade-off, showing the core-thermometer temperature differential (left y-axis, in blue), and the heater power required for core heating of $0.1$ K (right y-axis, in dark red), versus heater-thermometer spacing.
+Extended Data Figure 2: Measurement of temperature coefficient of resistance of the integrated Platinum thermometer.
 a, Experimental schematic depicting a test chip, containing a platinum resistor (referred to here as device under test (DUT)) fabricated on top of a SiN microresonator, placed on a temperature-stabilized hot plate chuck inside a vacuum chamber.
@@
 b, Measured electrical resistance of DUT versus temperature (dark-blue filled circles) plotted alongside a linear fit (magenta solid line).
-The integrated Platinum resistance thermometer, which allows the real-time measurement of the microresonator as described in the Main Text, is calibrated as shown in Fig.
-M2(a).
-A test device is placed on a temperature-controlled chuck using a thermally-conductive adhesive in order to ensure proper thermal contact.
-The setup is placed and operated inside a vacuum probe station in order to minimize unwanted convective cooling.
-Two probes, connected to a Keithley $2400$ series SourceMeter®, are brought in electrical contact with the two ends of the Pt resistor under test (labeled as device under test (DUT) in Fig.
-M2).
-At a given temperature of the chuck, the resistance $R$ of the DUT is measured by performing a current-voltage characterization.
-The voltage drop across the two contact pads is measured at various source currents, and a linear fit of the resulting data is obtained to extract $R$.
-This method is further explained in a prior manuscript [52].
-The source current values used are low enough ($<1$ mA) to avoid any self-heating of the resistor.
-This process is repeated for a range of chuck temperatures $T$.
-Figure M2(b) displays the results.
-Over the range of temperatures probed on this experiment, the resistance exhibits a linear trend.
-The temperature-coefficient of resistance is extracted by fitting the experimental data to Eq.
-1, and is determined to be $A=0.00172$.
-We note that DUT described here has a shorter fabricated length compared to the resistance thermometer described in the Main Text, leading to the lower base resistance.
+VI Data Availability
+The data shown in this paper are available in this Zenodo repository: https://doi.org/10.5281/zenodo.17144471.